OpenStudy (anonymous):

(sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2x

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (anonymous):

sinx (tanx cosx - cotx cosx) = 1 - 2 cos²x sinx ((sinx/cosx) cosx - (cosx/sinx) cosx) = ∙∙∙ sinx (sinx - cos²x/sinx) = ∙∙∙ sin²x - cos²x = ∙∙∙ (1 - cos²x) - cos²x = ∙∙∙ 1 - 2 cos²x = 1 - 2 cos²x